### v2

# Diffusion-Augmented Contrastive Learning: A Noise-Robust Encoder for Biosignal Representations

Abstract
Learning robust representations for biosignals is often hampered by the challenge of designing effective data augmentations. Traditional methods can fail to capture the complex variations inherent in physiological data. Within this context, we propose a novel hybrid framework, Diffusion-Augmented Contrastive Learning (DACL), that fuses concepts from diffusion models and supervised contrastive learning. The DACL framework operates on a latent space created by a lightweight Variational Autoencoder (VAE) trained on our novel Scattering Transformer (ST) features [12]. It utilizes the diffusion forward process as a principled data augmentation technique to generate multiple noisy ”views” of these latent embeddings. A U-Net style encoder is then trained with a supervised contrastive objective to learn a representation that balances class discrimination with robustness to noise across various diffusion timesteps. We evaluated this proof-of-concept method on the PhysioNet 2017 ECG dataset, achieving a competitive AUROC of 0.7815. This work establishes a new paradigm for representation learning by using the diffusion process itself to drive the contrastive objective, creating noise-invariant embeddings that demonstrate a strong foundation for class separability.
I Introduction
The efficacy of contrastive learning frameworks [1, 2, 13] fundamentally hinges upon the quality of data augmentations used to generate distinct yet semantically coherent ”views” of a data sample. For images, a rich set of heuristic augmentations is well-established. For physiological time-series, however, designing effective and meaning-preserving augmentations remains a significant and unsolved challenge, as simplistic transformations may not adequately capture the complex variations of the data manifold [3, 4, 15]. Simultaneously, diffusion models [14, 6] have emerged as powerful generative tools, exhibiting remarkable proficiency in learning fundamental data distributions through the modelling of a progressive noising and denoising process. The forward noising process provides a systematic approach for creating increasingly noisy iterations of a data point that stay close to the learned data manifold, a characteristic well utilized by denoising autoencoders [7]. Motivated by parallel advancements in diffusion models and contrastive learning, this paper addresses the challenge of creating effective data augmentations for physiological time-series signals. We introduce Diffusion-Augmented Contrastive Learning (DACL), a framework that replaces handcrafted augmentations with a principled, learned data augmentation strategy derived from the diffusion forward process. Instead of relying on heuristic transformations, DACL generates views by sampling embeddings at varying noise levels along the diffusion trajectory. This work operates within a latent space derived from our previously developed Scattering Transformer (ST) [12] features, leveraging their robust and structured nature. The key contributions of this paper are as follows:
  • Proposing DACL, a novel hybrid framework that effectively fuses diffusion models and supervised contrastive learning for noise-robust representation learning.
  • Introducing the use of the diffusion forward process as a principled, learned augmentation strategy, thereby eliminating the reliance on heuristic augmentations for biosignals.
  • Demonstrating that the proposed framework learns representations that are robust to noise while achieving competitive class discrimination on a standard biosignal benchmark.
The remainder of this paper is organized as follows: Section II reviews related work. The details of the proposed methodology is presented in Section III. Section IV presents the experimental setup and results, Section V discusses the findings, and Section VI concludes the paper.
II Related Work
Figure 1: Conceptual diagram of the DACL framework. Stage I involves training a VAE on all ST features to create a general-purpose latent space. In Stage II, this latent space is used for the main training loop. A triplet of latent vectors is sampled, and noisy views are generated via the diffusion forward process. These views are passed through a shared U-Net encoder, which is trained via a supervised triplet loss to produce noise-robust, class-discriminative embeddings.
II-A The Role of Augmentation in Contrastive Learning
Contrastive learning trains an encoder to be invariant to certain transformations while remaining sensitive to semantic differences [1, 2]. The choice of these transformations, or augmentations, is therefore of critical importance. While a well-established set of augmentations exists for images [8], the optimal augmentations for time-series are less clear. Common methods include jittering, scaling, magnitude warping, and adding noise [3], but these are often applied heuristically and may fail to maintain critical signal attributes. Research has shown that the performance of contrastive learning is highly sensitive to the chosen augmentation strategy [9], motivating a search for more principled and data-driven approaches. This need for robust representations is especially critical as new, powerful architectures for sequence modeling, such as state-space models, continue to emerge [17].
II-B Learning with Noise
Learning robust representations in the presence of noise is a long-standing objective in machine learning. Denoising Autoencoders (DAEs) [7] were an early and effective method, trained to reconstruct a clean input from a corrupted version. This objective forces the model to learn a compressed representation that captures the essential, low-dimensional features of the data manifold. More recently, diffusion models [6, 10] have formalized this concept into a powerful generative framework by learning to reverse a multi-step, controlled noising process. The intermediate steps of this process provide a rich spectrum of noisy-but-structured views of the data, which are inherently more meaningful than unstructured noise.
II-C Positioning of Our Work
To the best of our knowledge, our work is the first to propose the fusion of these two ideas for representation learning. We build on the established effectiveness of contrastive learning on structured features and use the powerful Scattering Transformer (ST) [12] as our feature foundation. Our core novelty lies in replacing handcrafted augmentations with views generated by the diffusion forward process. While other works have used generative models to create new data instances for contrastive learning [11], our approach is distinguished by its use of the generative process itself as the augmentation mechanism applied to existing samples. This is distinct from other recent approaches that leverage the pre-trained denoiser of a diffusion model as a feature extractor [18]; in contrast, we use the forward noising process itself to train a noise-robust encoder from scratch.
III Methodology
III-A Framework Overview
This paper introduces Diffusion-Augmented Contrastive Learning (DACL), a novel hybrid framework that synthesizes principles from diffusion models and supervised contrastive learning to learn noise-robust representations. The central tenet of the framework is the replacement of handcrafted data augmentations with a principled noising process derived from a diffusion model, as illustrated in Fig. 1. The framework operates in two distinct phases: first, learning a general-purpose latent space from ST features, and second, training a noise-robust encoder within that space using a diffusion-augmented contrastive objective.
III-B Foundational Feature Representation and Latent Space
The framework begins with fixed-size embeddings generated by our training-free Scattering Transformer (ST) [12]. To create a compressed and smooth manifold suitable for the diffusion process, a lightweight Variational Autoencoder (VAE) is trained on the ST features from the entire training dataset, including all available classes. Operating in a compressed latent space is a core principle of latent diffusion models, which have demonstrated high efficiency and quality in generative tasks [16]. This pre-training phase aims to allow the VAE to acquire a thorough representative latent space. The mean of the VAE’s learned posterior distribution, $\mu$, for each input feature is used as the clean latent vector $\mathbf{z}_{0}$ for the subsequent DACL training process.
III-C Core Mechanism: Diffusion-based Augmentation
The novelty of the proposed method lies in its data augmentation approach for the bio-signals. For a given clean latent vector $\mathbf{z}_{0}$, a noisy view at a randomly sampled timestep $t$ is generated using the standard diffusion forward process:
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\quad\text{where }\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ (1)
This process generates structured noise that follows the learned data manifold. This aims to create more meaningful augmentations than simply adding unstructured Gaussian noise.
III-D Mathematical Formulation
The central component of the learning phase is a U-Net style encoder, $\text{Enc}_{\theta}$, implemented here as a Multi-Layer Perceptron (MLP) suitable for operating on vector-based latent spaces. While not structurally a convolutional U-Net, we adopt this terminology to reflect its functional role as the conditional noise predictor, which is analogous to the U-Net’s role in seminal diffusion models. For a given triplet, views are generated as follows: an anchor and a positive view are created from the same initial latent vector ($\mathbf{z}_{0,A}=\mathbf{z}_{0,P}$) by applying the diffusion process with independently sampled noise and timesteps ($t_{A},t_{P}$). A negative view is generated from a latent vector of a different class, $\mathbf{z}_{0,N}$, at a random timestep $t_{N}$. These three noisy views ($\mathbf{z}_{t_{A},A}$, $\mathbf{z}_{t_{P},P}$, $\mathbf{z}_{t_{N},N}$) are then passed through the shared U-Net encoder to produce the final embeddings $\mathbf{e}_{A},\mathbf{e}_{P},\mathbf{e}_{N}$. The encoder is trained by minimizing a supervised triplet margin loss:
$L(\theta)=\mathbb{E}\left[\max(0,\|\mathbf{e}_{A}-\mathbf{e}_{P}\|_{2}^{2}-\|\mathbf{e}_{A}-\mathbf{e}_{N}\|_{2}^{2}+m)\right]$ (2)
This objective compels the encoder to learn representations that are invariant to the level of diffusion noise (by pulling $\mathbf{e}_{A}$ and $\mathbf{e}_{P}$ together) while simultaneously encouraging separability between classes (by pushing $\mathbf{e}_{A}$ and $\mathbf{e}_{N}$ apart).
III-E Training and Inference Procedure
The complete workflow, detailed in Algorithm 1, is a sequential two-phase process.
  • Training: First, the VAE is trained on all available ST features to learn the latent space representation. Second, the VAE encoder is frozen, and the U-Net encoder is trained exclusively on the contrastive task using triplets of latent vectors and their corresponding noisy views.
  • Inference: For a given test sample $\mathbf{x}_{test}$, it is passed through the frozen VAE encoder to obtain the corresponding clean latent vector $\mathbf{z}_{0,test}$. This vector is then passed through the trained U-Net encoder using a fixed, low-noise timestep (e.g., $t=1$) to produce a stable, noise-robust final embedding $\mathbf{e}_{test}$. A normal-class prototype, $\boldsymbol{\mu}_{normal}$, is pre-computed by taking the mean of the final embeddings of all normal samples from the training set. The anomaly score is determined by the Euclidean distance $\|\mathbf{e}_{test}-\boldsymbol{\mu}_{normal}\|_{2}$.
Algorithm 1 Training and Inference for DACL
1:Require: Training ST features $\mathcal{X}=\{\mathbf{x}_{i}\}_{i=1}^{N}$, Labels $\mathcal{Y}=\{y_{i}\}_{i=1}^{N}$.
2:Ensure: A trained DACL encoder $\text{Enc}_{\theta}$ and a VAE encoder $E_{VAE}$.
3:
4:// Phase 1: Train VAE and obtain latent vectors
5:Train VAE on all features $\mathcal{X}$ to get encoder $E_{VAE}$.
6:Compute clean latent vectors $\mathcal{Z}_{0}\leftarrow E_{VAE}(\mathcal{X})$.
7:
8:// Phase 2: Train DACL Encoder
9:Initialize U-Net encoder parameters $\theta$.
10:for each training epoch do
11:  for each training step do
12:   Sample anchor latent vector $\mathbf{z}_{0,A}$ and negative $\mathbf{z}_{0,N}$.
13:   Set positive latent vector $\mathbf{z}_{0,P}=\mathbf{z}_{0,A}$.
14:// Generate noisy views via diffusion forward process
15:   Sample timesteps $t_{A},t_{P},t_{N}$ and noise $\boldsymbol{\epsilon}_{A},\boldsymbol{\epsilon}_{P},\boldsymbol{\epsilon}_{N}$.
16:   $\mathbf{z}_{t_{A},A}\leftarrow\sqrt{\bar{\alpha}_{t_{A}}}\mathbf{z}_{0,A}+\sqrt{1-\bar{\alpha}_{t_{A}}}\boldsymbol{\epsilon}_{A}$
17:   $\mathbf{z}_{t_{P},P}\leftarrow\sqrt{\bar{\alpha}_{t_{P}}}\mathbf{z}_{0,P}+\sqrt{1-\bar{\alpha}_{t_{P}}}\boldsymbol{\epsilon}_{P}$
18:   $\mathbf{z}_{t_{N},N}\leftarrow\sqrt{\bar{\alpha}_{t_{N}}}\mathbf{z}_{0,N}+\sqrt{1-\bar{\alpha}_{t_{N}}}\boldsymbol{\epsilon}_{N}$
19:// Get final embeddings from U-Net encoder
20:   $\mathbf{e}_{A}\leftarrow\text{Enc}_{\theta}(\mathbf{z}_{t_{A},A},t_{A})$
21:   $\mathbf{e}_{P}\leftarrow\text{Enc}_{\theta}(\mathbf{z}_{t_{P},P},t_{P})$
22:   $\mathbf{e}_{N}\leftarrow\text{Enc}_{\theta}(\mathbf{z}_{t_{N},N},t_{N})$
23:   Compute triplet loss $L_{triplet}$ on $(\mathbf{e}_{A},\mathbf{e}_{P},\mathbf{e}_{N})$.
24:   Update $\theta$ using gradient descent on $L_{triplet}$.
25:  end for
26:end for
27:
28:// Inference Procedure
29:Input: A test feature vector $\mathbf{x}_{test}$.
[…]
32:Compute normal class prototype $\boldsymbol{\mu}_{normal}=\text{mean}(\text{Enc}_{\theta}(\mathcal{Z}_{0,normal},t=1))$.
33:return Anomaly Score $\leftarrow\|\mathbf{e}_{test}-\boldsymbol{\mu}_{normal}\|_{2}$
IV Experiments and Results
IV-A Dataset and Preprocessing
We evaluated our proposed method on the PhysioNet 2017 Challenge Dataset [5]. This provides a benchmark for ECG classification-related tasks. It contains recordings of Normal, Atrial Fibrillation (AF), and Other rhythms. For this study, the problem was framed as a binary anomaly detection task. The ’Normal’ recordings served as the normal class, while the ’Atrial Fibrillation’ and ’Other’ rhythm recordings were combined to form a single ’Anomaly’ class. The dataset was partitioned by patient to ensure subject independence, with 70 percent allocated for training and 30 percent for testing.
IV-B Evaluation Protocol and Baselines
Features for all experiments were extracted using our previously developed Scattering Transformer (ST) [12]. The quality of the learned representations was evaluated by training a simple linear classifier on the frozen embeddings produced by the final encoder. Performance was measured by the patient-level Area Under the Receiver Operating Characteristic curve (AUROC). We compared our model against two conceptual baselines:
  1. 1.
    Contrastive + Gaussian Aug. (Baseline 1): A standard supervised contrastive model where augmentations are created by adding simple, handcrafted Gaussian noise to the latent vectors.
  2. 2.
    Denoising Autoencoder (Baseline 2): A standard DAE trained to reconstruct clean latent embeddings from noisy versions generated via the same diffusion forward process.
These baselines were specifically chosen to isolate the contributions of our method. Baseline 1 tests the benefit of structured, manifold-aware diffusion noise over unstructured Gaussian noise. Baseline 2 tests the benefit of a discriminative contrastive objective over a purely reconstructive objective when using the same noise generation process.
IV-C Results
All models were trained using the Adam optimizer. The main DACL framework used a VAE with a 32-dimensional latent space and a U-Net encoder trained for 200 epochs with a learning rate of 1e-4 and a batch size of 64.
Our proposed DACL framework achieved a competitive patient-level AUROC of 0.7815 on the PhysioNet 2017 dataset. While this proof-of-concept result does not outperform heavily tuned, fully supervised classifiers, its performance is noteworthy, particularly for a model learning representations via such a novel paradigm. The primary comparison of interest is against the conceptual baselines designed to isolate the benefits of our approach, as shown in Table I. The corresponding patient-level ROC curve is shown in Fig. 2.
TABLE I: Performance comparison of the DACL framework against conceptual baselines on PhysioNet 2017.
Model AUROC
DACL (Ours) 0.7815
Contrastive + Gaussian Aug. 0.6716
Denoising Autoencoder 0.7532
IV-D Ablation Study: Impact of Diffusion Timestep
To isolate and quantify the impact of the diffusion timestep $t$ on the quality of the learned representation, we performed an ablation study. We trained separate DACL encoders using views generated exclusively from distinct ranges of $t$: ’Early’ (1-16), ’Mid’ (17-33), and ’Late’ (34-50). As shown in Fig. 4, there is a clear positive correlation between the noise level of the augmentations and the final model performance. Using only ’Early’ timesteps resulted in the lowest AUROC, whereas using only ’Late’ timesteps achieved the highest performance. This implies that challenging the encoder with highly corrupted views forces it to learn a more robust and discriminative representation by focusing only on the most essential, noise-invariant features.
Figure 2: Patient-Level ROC Curve for the DACL model on PhysioNet 2017 (AUROC = 0.7815).
Figure 3: t-SNE visualization of the test embeddings generated by the DACL encoder.
Figure 4: Results from the diffusion timestep ablation study. The plot shows patient-level AUROC as a function of the timestep range used for generating contrastive views. Performance clearly improves with higher noise levels (’Late’ timesteps), demonstrating that more challenging augmentations force the model to learn a more robust representation.
V Discussion
This study introduced DACL, a proof-of-concept framework that successfully demonstrates the viability of fusing diffusion and contrastive learning. The competitive AUROC of 0.7815 validates the use of the diffusion forward process as a principled augmentation strategy. This approach not only improves upon the classic Denoising Autoencoder but also yields a significant 11-point AUROC gain over a standard contrastive framework using simple Gaussian noise augmentation. This result confirms the relative advantage of the learned, manifold-aware noise generated by the diffusion process compared to heuristic augmentations. The t-SNE visualization of the resulting embedding space is given in Fig. 3. The visualization reveals a nascent class structure, though with significant overlap still present, which is consistent with the achieved AUROC. Furthermore, our ablation study on the diffusion timestep (Fig. 4) provided a key insight: the most effective representations were learned when the contrastive objective was driven by high-noise augmentations from late in the diffusion process. This confirms the model’s ability to learn truly noise-invariant features when presented with sufficiently challenging augmentations. The primary implication of this work is methodological. It opens a new avenue for research into learned, generative augmentations for contrastive and self-supervised learning. This could be particularly valuable in domains like biosignals, where defining effective handcrafted augmentations is a major unresolved challenge. While the results are promising, it is important to acknowledge the limitations of this proof-of-concept study. The performance did not match that of fully supervised, state-of-the-art methods. Nevertheless, the results suggest that feasibility of the idea and that further optimization of the U-Net encoder architecture and the VAE latent space is required to achieve top-tier performance.
VI Conclusion
This work introduced Diffusion-Augmented Contrastive Learning (DACL), a novel framework that successfully synthesizes diffusion models and supervised contrastive learning. We proposed using the diffusion forward process as a principled and learned data augmentation technique to generate the ”views” required for a contrastive learning objective. Our assessment of the PhysioNet 2017 dataset confirmed the efficacy of our method, attaining a competitive AUROC of 0.7815. This methodological study, although not surpassing fully supervised approaches, successfully developed a novel paradigm for learning noise-invariant representations that underpin class discrimination. This study aims to pave the way for further research in noise robust representation learning for biosignals and other fields where augmentations are complex.